stat question 10

given independent random variables with means and standarddeviations as shown, find the mean and standard deviation of eachof these variables.
MeanStandard Deviation
X8012
Y123

a) X-20b).5Yc)X+Yd)X-Ye)Y1+Y2
Answer #1
Given that
E(X) = 80
E(Y) = 12
S.D(X) = 12
S.D(Y) 3

a) Mean
E(X-20) = E(X) -20
= 80-20 =60
Standard deviation
S.D(X-20)=√V(X-20) =√V(X)[since V(constant) = 0]
=S.D(X) = 12
b) Mean
E(5Y) = 5E(Y)
=5(12) = 60
S.D(5Y) = √V(5y) = √52 V(Y) = 5S.D(Y) = 5(3)= 15
[SinceV(constant with variable) = (constant)2V(variable)]

Similary you can calculate the remaining bits
Hope this will help you
answered by: morgan2